# Safety with Agency: Human-Centered Safety Filter with Application to AI-Assisted Motorsports

### v3

@@ I Introduction
Recent developments in robot safety provide an exciting opportunity for enhancing human safety and performance in high-stakes situations. However, augmenting human decision-making with artificial intelligence (AI) in a trustworthy way remains an open problem. A human–AI team in a performance car racing [1, 2, 3] is a representative, yet challenging example, as it pushes safety to the limit. How should the AI co-pilot assist the human without dulling the driver’s competitive edge? Should the AI discourage a human from attempting a risky overtaking maneuver on a sharp turn? When an AI system assists humans in such safety-critical and time-sensitive tasks, maintaining human agency is critical. We need to ensure human awareness of the AI system’s current intent and operating mode, thus avoiding the notorious and sometimes fatal “automation surprise” [4, 5].
Safety filters [6, 7] have become an effective approach to ensure safety under an operational design domain (ODD), i.e., a clearly defined set of operating conditions for robots to work properly and safely [8, 9]. Safety filters have been deployed on a wide range of autonomous systems, such as automated vehicles [10, 11], legged robots [12, 13, 14], and aerial navigation [15, 16]. Traditional model-based numerical approaches for safety filter synthesis [17] result in safety guarantees by design, but they are unable to scale up due to the “curse of dimensionality” [18]. Recent research focuses on neural approximation of safety filters [19, 20, 21, 22, 23] that can scale to tens [14, 24] and even hundreds of state variables [25]. While existing safety filters effectively maintain safety, their use in human–AI shared autonomy can result in abrupt, discontinuous interventions that disregard the human operator’s intentions. This undermines the driver’s sense of being in control, creating an uncomfortable and unenjoyable experience. Moreover, such unpredictable, non-transparent behavior can erode confidence and trust in the AI assistant, ultimately degrading team performance and causing the human driver to lose their strategic edge.
@@ II-A Human-Interactive Safety Filters
A safety filter is a supervisory control scheme that continuously monitors the operation of an autonomous system and intervenes, when necessary, by adjusting its planned actions to prevent potential catastrophic failures. Safety filters have been increasingly used in high-stakes autonomy applications, ranging from autonomous driving [26, 11, 27, 28], to aerial navigation [29, 30, 20, 16, 31], and to legged locomotion [12, 24, 14, 23]. Recent work by Hsu et al. [7] provides a unified analysis framework for various safety filters, including Hamilton–Jacobi (HJ) reachability [17, 18, 16], control barrier functions [32, 20, 33], model predictive control [34, 35], and Lyapunov methods [36]. In general, synthesis of safety filters can be computationally challenging, especially for systems with high-dimensional state space and complex dynamics. Deep learning has proven to effectively scale up the computation of safety controllers [19, 37, 25, 14, 23]. More recently, methods have been developed that treat these learned neural controllers as an untrusted fallback within a safety filter framework, and robust safety guarantees can be subsequently obtained through runtime verification algorithms such as convex optimization [38, 39, 40], forward reachable sets rollouts [22], and conformal prediction [41].
@@
When robots are deployed around humans, ensuring safety is paramount to enable their trustworthy integration into people’s everyday lives. However, enforcing safety becomes particularly challenging in human-interactive settings due to coupled motion, limited communication, and potentially conflicting objectives between robots and their human peers. Early attempts at safe human–robot interaction focus on achieving robust safety by safeguarding against worst-case human decisions [42, 26, 43], which may lead to overly conservative robot behaviors [44]. Recent research effort has been devoted to designing safety filters that adapt to human decision-making, in hope of improving the robot’s task performance without compromising safety. One popular approach is filter-aware motion planning, which incorporates predictions of the safety filter’s behaviors into the robot’s task policy [26, 45]. This strategy allows the robot to avoid abrupt safety overrides by preempting future costly interventions triggered by unlikely human actions. Another line of research aims at reducing conservativeness by dynamically adjusting the safety filter’s ODD according to the robot’s evolving uncertainty about the human [46, 25, 47].
@@
While existing human-interactive safety filters enable robots to interact safely and efficiently with other humans, similar formulations in human–robot shared control settings remain scarce. Recently, research efforts have focused on preserving human agency while enhancing safety in shared autonomy [48, 49, 50]. However, these approaches do not define a clear ODD and lack the principled safety analysis that a safety filter provides, often leading to elevated failure rates. Moreover, some methods rely on knowing the human operator’s policy a priori, limiting their robustness when working with groups of human operators who have diverse intentions and skill levels [48, 50].
@@
In this work, we draw inspiration from filter-aware planning to design a safety filter that minimally modifies human actions. Our proposed HCSF preserves the principled safety analysis inherited from safety filter theory—in particular, from HJ reachability and control barrier functions—while avoiding any explicit model representation of human intentions.
@@ II-B AI-Enabled Motorsports
While modern AI systems surpass human intelligence in competitive sports [51, 52], their potential to augment human decision-making is underexplored. High-speed performance car racing presents a domain where safety and seamless collaboration are required to enable a competitive human–AI team—the AI co-pilot must assist the human without dulling the driver’s competitive edge.
Wurman et al. [1] demonstrate for the first time that a well-trained neural policy can win a head-to-head competition against some of the world’s best drivers in a car racing game. Follow-up works further improve the AI competitiveness via reasoning strategic interactions with data-driven modeling of opponent behaviors [53] and blending model-based dynamic game strategy with data-driven prior knowledge [54, 55]. Comparing to fully automated AI motorsports, human–AI collaborative car racing is an emerging, yet relatively underexplored research area. Gopinath et al. [3] propose a multi-task imitation learning approach that enables an automated coaching system that interacts with the student similar to a human teacher. DeCastro et al. [2] enhance the performance of human–AI teams in car racing by learning a policy that infers and aligns with human intents leveraging a world model. While AI agents in motorsports have shown promising performance, ensuring the safety of human drivers remains largely unaddressed in a principled manner. Chen et al. [56] present preliminary results on approximate learning-based safety analysis for autonomous racing, but their approach is limited to the single-car, fully automated setting.
@@ III Preliminaries and Problem Formulation
:-${\mathcal{F}}\coloneq\{{x}\in{\mathcal{X}}\mid g({x})<0\}.$ (2)
@@ III-A Safety Filters
A safety filter [7] is an automated process that continuously monitors the system and intervenes, if deemed necessary, by modifying a candidate action given by the task policy ${{\pi}^{\text{task}}}$ to prevent a potentially catastrophic safety failure in the future. Specifically, instead of directly applying the task action ${u}_{t}={{\pi}^{\text{task}}}({x}_{t})$, the robot uses an action based on safety filtering:
@@
The specific function form of ${\phi}$ depends on the intervention type of a safety filter, which includes, e.g., switching, transition, and optimization [7, Sec. 3]. The safety filter only prevents the use of a task action that would compromise future safety, allowing the robot to maintain safety without needing to modify its entire behavior. In this paper, we use HJ reachability analysis to synthesize safety filter ${\phi}$ [18, 17].
@@ III-B Hamilton-Jacobi Reachability Analysis
We aim to design a safety filter which, given ODD elements ${\mathcal{F}}$ and ${\mathcal{U}}$, keeps the robot within the maximal safe set ${\mathcal{S}}^{*}\subset{\mathcal{F}}^{\mathsf{c}}\subset{\mathcal{X}}$, where ${\mathcal{F}}^{\mathsf{c}}={\mathcal{X}}-{\mathcal{F}}$. This set ${\mathcal{S}}^{*}$ consists of all states from which there exists a control policy that indefinitely prevents the robot from entering ${\mathcal{F}}$. In theory, ${\mathcal{S}}^{*}$ can be computed using Hamilton-Jacobi (HJ) reachability analysis, which employs level-set methods to recast the safety filter synthesis problem as an optimal control problem. Its solution follows from solving the dynamic programming safety Bellman equation [17]:
@@
For subsequent extension to our proposed HCSF, we adopt the ${\mathcal{Q}}$-function [57]—a notion widely used in reinforcement learning (RL)—to modify (4) into the state–action safety Bellman equation:
@@
Prior work [16, 17, 18, 19, 29] has established LRSF as a fundamental framework applicable for all HJ reachability analysis-based safety filters. This is due to its straightforward yet effective design for enforcing safety and its “least-restrictive” nature, allowing the task-oriented policy full freedom until the system reaches the boundary of the maximal safe set.
However, LRSF also has notable drawbacks. First, its task-agnostic fallback policy often leads to discontinuous and jerky interventions [7]. This issue can become more pronounced in a shared autonomy setting, where human operators might feel surprised and confused by abrupt overrides [4, 5]. Even though LRSF offers maximum freedom until the system reaches the boundary of ${\mathcal{S}}^{*}$, its interventions that do not take into account the human’s input risk diminishing the operator’s sense of control. Furthermore, the time and space complexities required to solve the Bellman equation (4) scale exponentially with the state-space dimension, rendering grid-based dynamic programming infeasible for real-world safe robot control.
Our proposed HCSF addresses these limitations by synthesizing an output that minimally deviates from the human operator’s input, thereby enhancing both agency and smoothness while still enforcing safety (Section IV). In addition, in Section V, we leverage recent advances in safety RL [19, 37, 23] to approximate the ${\mathcal{Q}}$-function via RL, enabling the synthesis of a best-effort fallback policy $\faShield*{{\pi}^{\text{\tiny{\faShield*}}}}(\cdot)$ for high-dimensional systems.
@@ III-C Discrete-Time Control Barrier Functions
Definition 1 (Discrete-time CBF [13]).
@@
An HJ safety value function ${V}(\cdot)$ is closely linked to a DCBF in the sense that, if ${V}(\cdot)$ is continuously differentiable, it automatically qualifies as a valid CBF for the maximal safe set ${\mathcal{S}}^{*}$ [7, Sec. 3.2]. This insight enables the use of ${V}(\cdot)$ in a smooth CBF safety filter rather than the LRSF alternative—an approach that underpins our proposed HCSF.
@@
Definition 1 could be relaxed such that (8) is required to hold for all ${x}_{t}\in\{{x}\in{\mathcal{X}}\mid h({x})\geq 0\}$. A control input ${u}_{t}\in{\mathcal{U}}$ that satisfies (9b) for any function $h(\cdot)$ meeting the relaxed DCBF definition still renders the 0-superlevel set of $h(\cdot)$ forward invariant. Such relaxation of (8) still guarantees safety, but loses the set attractiveness property for the 0-superlevel set of $h(\cdot)$ [58].
@@ IV Smooth Human-Centered Safety Filter for Shared Autonomy
Conventional CBF safety filter methods similar to (9) typically require knowledge of the system’s dynamics [32, 10, 13, 15, 58, 31], even when using learned barrier functions [20, 33, 59, 60, 61, 62]. This requirement arises for one or both of the following reasons: 1) either a full-order or a simplified dynamical model of the system is utilized to synthesize CBF candidates, and 2) knowledge of the dynamics (e.g., control affine model) is leveraged at runtime to enforce the CBF safety constraint within an optimal control problem (OCP) (e.g., quadratic program).
@@
While some works explicitly aimed to build and deploy model-free CBF safety filters, they have so far fallen short of being fully model-free—relying on some combination of simplified models of the system dynamics [63, 64, 65], predefined low-level controllers [64, 65], and handcrafted fallback policies (e.g., evading maneuvers) [66]. Additionally, recent efforts in learning a CBF for latent state representations have proven to be effective for partially observable systems, but they still require a control affine dynamical model [67]. Such reliance on knowledge of the system dynamics and the deployment environment can significantly limit the applicability of CBF s in complex, real-world scenarios where the dynamical model is often unknown and should be treated as a black-box. On the other hand, a model-free algorithm for learning a policy together with a barrier certificate was proposed recently [68], but it cannot be used to build a safety filter because the learned policy must be deployed at all times. Finally, we acknowledge a preprint reporting concurrent efforts toward a model-free state–action CBF safety filter [69]. However, it addresses a finite-horizon safety problem and enforces safety at runtime like a smooth least restrictive safety filter [70] rather than a CBF one, fundamentally differing from our work in both mathematical formulation and enforcement of safety.
@@
To this end, we introduce, to the best of our knowledge, the first CBF safety filter that is fully model-free. We first show that the safety value function ${V}(\cdot)$ is itself a valid DCBF in the sense of Definition 1. Then, using the state–action safety value function ${\mathcal{Q}}(\cdot,\cdot)$ which could be learned scalably through black-box interactions with the system via model-free RL-based HJ reachability analysis (Section V-A), we propose a method of enforcing a novel Q-CBF safety constraint that does not require any information regarding the dynamics. While theoretical safety guarantees are contingent on the validity of the learned CBF (which may be established through statistical analysis [41] or model-based verification [20, 38]), this is not the focus of our research. Instead, we show our safety filter achieves an extremely high empirical safe rate and effectively preserves human agency.
@@
The optimization problem in Eq. (11) is recursively feasible for $\forall\gamma\in[0,1)$, given any initial state ${x}\in{\mathcal{S}}$.

### v4

@@ I Introduction
Recent developments in robot safety provide an exciting opportunity for enhancing human safety and performance in high-stakes situations. However, augmenting human decision-making with artificial intelligence (AI) in a trustworthy way remains an open problem. A human–AI team in a performance car racing [wurman2022outracing, decastro2024dreaming, gopinath2024computational] is a representative, yet challenging example, as it pushes safety to the limit. How should the AI co-pilot assist the human without dulling the driver’s competitive edge? Should the AI discourage a human from attempting a risky overtaking maneuver on a sharp turn? When an AI system assists humans in such safety-critical and time-sensitive tasks, maintaining human agency is critical. We need to ensure human awareness of the AI system’s current intent and operating mode, thus avoiding the notorious and sometimes fatal “automation surprise” [sarter1997team, jamieson2022b737].
Safety filters [bastani2021safe, hsu2023safety] have become an effective approach to ensure safety under an operational design domain (ODD), i.e., a clearly defined set of operating conditions for robots to work properly and safely [sae3259, RR-2662]. Safety filters have been deployed on a wide range of autonomous systems, such as automated vehicles [zeng2021safety, tearle2021predictive], legged robots [hsu2015control, agrawal2017discrete, nguyen2024gameplay], and aerial navigation [singletary2022onboard, chen2021fastrack]. Traditional model-based numerical approaches for safety filter synthesis [mitchell2005time] result in safety guarantees by design, but they are unable to scale up due to the “curse of dimensionality” [bansal2017hamilton]. Recent research focuses on neural approximation of safety filters [fisac2019bridging, robey2020learning, bansal2021deepreach, hsu2023isaacs, wang2024magics] that can scale to tens [nguyen2024gameplay, he2024agile] and even hundreds of state variables [hu2023deception]. While existing safety filters effectively maintain safety, their use in human–AI shared autonomy can result in abrupt, discontinuous interventions that disregard the human operator’s intentions. This undermines the driver’s sense of being in control, creating an uncomfortable and unenjoyable experience. Moreover, such unpredictable, non-transparent behavior can erode confidence and trust in the AI assistant, ultimately degrading team performance and causing the human driver to lose their strategic edge.
@@ II-A Human-Interactive Safety Filters
A safety filter is a supervisory control scheme that continuously monitors the operation of an autonomous system and intervenes, when necessary, by adjusting its planned actions to prevent potential catastrophic failures. Safety filters have been increasingly used in high-stakes autonomy applications, ranging from autonomous driving [leung2020infusing, tearle2021predictive, hu2022sharp, brat2023autonomy], to aerial navigation [chen2018hamilton, squires2018constructive, robey2020learning, chen2021fastrack, oh2023safety], and to legged locomotion [hsu2015control, he2024agile, nguyen2024gameplay, wang2024magics]. Recent work by Hsu et al. [hsu2023safety] provides a unified analysis framework for various safety filters, including Hamilton–Jacobi (HJ) reachability [mitchell2005time, bansal2017hamilton, chen2021fastrack], control barrier functions [ames2016control, robey2020learning, lindemann2021learning], model predictive control [li2020safe, wabersich2021predictive], and Lyapunov methods [chow2018lyapunov]. In general, synthesis of safety filters can be computationally challenging, especially for systems with high-dimensional state space and complex dynamics. Deep learning has proven to effectively scale up the computation of safety controllers [fisac2019bridging, hsu2021safety, hu2023deception, nguyen2024gameplay, wang2024magics]. More recently, methods have been developed that treat these learned neural controllers as an untrusted fallback within a safety filter framework, and robust safety guarantees can be subsequently obtained through runtime verification algorithms such as convex optimization [hu2020reach, everett2021reachability, gates2023scalable], forward reachable sets rollouts [hsu2023isaacs], and conformal prediction [lin2024verification].
@@
When robots are deployed around humans, ensuring safety is paramount to enable their trustworthy integration into people’s everyday lives. However, enforcing safety becomes particularly challenging in human-interactive settings due to coupled motion, limited communication, and potentially conflicting objectives between robots and their human peers. Early attempts at safe human–robot interaction focus on achieving robust safety by safeguarding against worst-case human decisions [fisac2018probabilistically, leung2020infusing, tian2022safety], which may lead to overly conservative robot behaviors [trautman2010unfreezing]. Recent research effort has been devoted to designing safety filters that adapt to human decision-making, in hope of improving the robot’s task performance without compromising safety. One popular approach is filter-aware motion planning, which incorporates predictions of the safety filter’s behaviors into the robot’s task policy [leung2020infusing, hu2024active]. This strategy allows the robot to avoid abrupt safety overrides by preempting future costly interventions triggered by unlikely human actions. Another line of research aims at reducing conservativeness by dynamically adjusting the safety filter’s ODD according to the robot’s evolving uncertainty about the human [bajcsy2021analyzing, hu2023deception, pandya2024robots].
@@
While existing human-interactive safety filters enable robots to interact safely and efficiently with other humans, similar formulations in human–robot shared control settings remain scarce. Recently, research efforts have focused on preserving human agency while enhancing safety in shared autonomy [broad2019highly, schaff2020residual, reddy2018shared]. However, these approaches do not define a clear ODD and lack the principled safety analysis that a safety filter provides, often leading to elevated failure rates. Moreover, some methods rely on knowing the human operator’s policy a priori, limiting their robustness when working with groups of human operators who have diverse intentions and skill levels [broad2019highly, reddy2018shared].
@@
In this work, we draw inspiration from filter‐aware planning to design a safety filter that minimally modifies human actions. Our proposed HCSF preserves the principled safety analysis inherited from safety filter theory—in particular, from HJ reachability and control barrier functions—while avoiding any explicit model representation of human intentions.
@@ II-B AI-Enabled Motorsports
While modern AI systems surpass human intelligence in competitive sports [mnih2015human, kaufmann2023champion], their potential to augment human decision-making is underexplored. High-speed performance car racing presents a domain where safety and seamless collaboration are required to enable a competitive human–AI team—the AI co-pilot must assist the human without dulling the driver’s competitive edge.
Wurman et al. [wurman2022outracing] demonstrate for the first time that a well-trained neural policy can win a head-to-head competition against some of the world’s best drivers in a car racing game. Follow-up works further improve the AI competitiveness via reasoning strategic interactions with data-driven modeling of opponent behaviors [chen2023learn] and blending model-based dynamic game strategy with data-driven prior knowledge [lidard2024blending, hu2024think]. Comparing to fully automated AI motorsports, human–AI collaborative car racing is an emerging, yet relatively underexplored research area. Gopinath et al. [gopinath2024computational] propose a multi-task imitation learning approach that enables an automated coaching system that interacts with the student similar to a human teacher. DeCastro et al. [decastro2024dreaming] enhance the performance of human–AI teams in car racing by learning a policy that infers and aligns with human intents leveraging a world model. While AI agents in motorsports have shown promising performance, ensuring the safety of human drivers remains largely unaddressed in a principled manner. Chen et al. [chen2021safe] present preliminary results on approximate learning-based safety analysis for autonomous racing, but their approach is limited to the single-car, fully automated setting.
@@ III Preliminaries and Problem Formulation
${\mathcal{F}}\coloneq\{{x}\in{\mathcal{X}}\mid g({x})<0\}.$ (2)
@@ III-A Safety Filters
A safety filter [hsu2023safety] is an automated process that continuously monitors the system and intervenes, if deemed necessary, by modifying a candidate action given by the task policy ${\pi}^{\text{task}}$} to prevent a potentially catastrophic safety failure in the future. Specifically, instead of directly applying the task action ${u}_{t}={{\pi}^{\text{task}}}({x}_{t})$, the robot uses an action based on safety filtering:
@@
The specific function form of $\phi$} depends on the intervention type of a safety filter, which includes, e.g., switching, transition, and optimization [hsu2023safety, Sec. 3]. The safety filter only prevents the use of a task action that would compromise future safety, allowing the robot to maintain safety without needing to modify its entire behavior. In this paper, we use HJ reachability analysis to synthesize safety filter $\phi$} [bansal2017hamilton, mitchell2005time].
@@ III-B Hamilton-Jacobi Reachability Analysis
We aim to design a safety filter which, given ODD elements $\mathcal{F}$} and $\mathcal{U}$}, keeps the robot within the maximal safe set ${\mathcal{S}}^{*}\subset{\mathcal{F}}^{\mathsf{c}}\subset{\mathcal{X}}$, where ${\mathcal{F}}^{\mathsf{c}}={\mathcal{X}}$-$\mathcal{F}$}. This set $\mathcal{S}$$^{*}$ consists of all states from which there exists a control policy that indefinitely prevents the robot from entering $\mathcal{F}$}. In theory, $\mathcal{S}$$^{*}$ can be computed using Hamilton-Jacobi (HJ) reachability analysis, which employs level-set methods to recast the safety filter synthesis problem as an optimal control problem. Its solution follows from solving the dynamic programming safety Bellman equation [mitchell2005time]:
@@
For subsequent extension to our proposed HCSF, we adopt the $\mathcal{Q}$}-function [watkins1992q]—a notion widely used in reinforcement learning (RL)—to modify (4) into the state–action safety Bellman equation:
@@
Prior work [chen2021fastrack, mitchell2005time, bansal2017hamilton, fisac2019bridging, chen2018hamilton] has established LRSF as a fundamental framework applicable for all HJ reachability analysis-based safety filters. This is due to its straightforward yet effective design for enforcing safety and its “least-restrictive” nature, allowing the task-oriented policy full freedom until the system reaches the boundary of the maximal safe set.
However, LRSF also has notable drawbacks. First, its task-agnostic fallback policy often leads to discontinuous and jerky interventions [hsu2023safety]. This issue can become more pronounced in a shared autonomy setting, where human operators might feel surprised and confused by abrupt overrides [sarter1997team, jamieson2022b737]. Even though LRSF offers maximum freedom until the system reaches the boundary of $\mathcal{S}$$^{*}$, its interventions that do not take into account the human’s input risk diminishing the operator’s sense of control. Furthermore, the time and space complexities required to solve the Bellman equation (4) scale exponentially with the state-space dimension, rendering grid-based dynamic programming infeasible for real-world safe robot control.
[…]
@@ III-C Discrete-Time Control Barrier Functions
Definition 1 (Discrete-time CBF [agrawal2017discrete]).
@@
An HJ safety value function ${V}(\cdot)$ is closely linked to a DCBF in the sense that, if ${V}(\cdot)$ is continuously differentiable, it automatically qualifies as a valid CBF for the maximal safe set $\mathcal{S}$$^{*}$ [hsu2023safety, Sec. 3.2]. This insight enables the use of ${V}(\cdot)$ in a smooth CBF safety filter rather than the LRSF alternative—an approach that underpins our proposed HCSF.
@@
Definition 1 could be relaxed such that (8) is required to hold for all ${x}_{t}\in\{{x}\in{\mathcal{X}}\mid h({x})\geq 0\}$. A control input ${u}_{t}\in{\mathcal{U}}$ that satisfies (9b) for any function $h(\cdot)$ meeting the relaxed DCBF definition still renders the 0-superlevel set of $h(\cdot)$ forward invariant. Such relaxation of (8) still guarantees safety, but loses the set attractiveness property for the 0-superlevel set of $h(\cdot)$ [cortez2021robust].
@@ IV Smooth Human-Centered Safety Filter for Shared Autonomy
Conventional CBF safety filter methods similar to (9) typically require knowledge of the system’s dynamics [ames2016control, zeng2021safety, agrawal2017discrete, singletary2022onboard, cortez2021robust, oh2023safety], even when using learned barrier functions [robey2020learning, lindemann2021learning, dawson2022safe, so2024train, zhang2025gcbf+, xiao2023barriernet]. This requirement arises for one or both of the following reasons: 1) either a full-order or a simplified dynamical model of the system is utilized to synthesize CBF candidates, and 2) knowledge of the dynamics (e.g., control affine model) is leveraged at runtime to enforce the CBF safety constraint within an optimal control problem (OCP) (e.g., quadratic program).
@@
While some works explicitly aimed to build and deploy model-free CBF safety filters, they have so far fallen short of being fully model-free—relying on some combination of simplified models of the system dynamics [li2022bridging, molnar2021model, cohen2024safety], predefined low-level controllers [molnar2021model, cohen2024safety], and handcrafted fallback policies (e.g., evading maneuvers) [squires2021model]. Additionally, recent efforts in learning a CBF for latent state representations have proven to be effective for partially observable systems, but they still require a control affine dynamical model [kumar2024latentcbf]. Such reliance on knowledge of the system dynamics and the deployment environment can significantly limit the applicability of CBF s in complex, real-world scenarios where the dynamical model is often unknown and should be treated as a black-box. On the other hand, a model-free algorithm for learning a policy together with a barrier certificate was proposed recently [yang2023model], but it cannot be used to build a safety filter because the learned policy must be deployed at all times. Finally, we acknowledge a preprint reporting concurrent efforts toward a model-free state–action CBF safety filter [he2023state]. However, it addresses a finite-horizon safety problem and enforces safety at runtime like a smooth least restrictive safety filter [borquez2024safety] rather than a CBF one, fundamentally differing from our work in both mathematical formulation and enforcement of safety.
@@
To this end, we introduce, to the best of our knowledge, the first CBF safety filter that is fully model-free. We first show that the safety value function ${V}(\cdot)$ is itself a valid DCBF in the sense of Definition 1. Then, using the state–action safety value function ${\mathcal{Q}}(\cdot,\cdot)$ which could be learned scalably through black-box interactions with the system via model-free RL-based HJ reachability analysis (Section V-A), we propose a method of enforcing a novel Q-CBF safety constraint that does not require any information regarding the dynamics. While theoretical safety guarantees are contingent on the validity of the learned CBF (which may be established through statistical analysis [lin2024verification] or model-based verification [robey2020learning, hu2020reach]), this is not the focus of our research. Instead, we show our safety filter achieves an extremely high empirical safe rate and effectively preserves human agency.
@@
The optimization problem in Eq. (11) is recursively feasible for $\forall\gamma\in[0,1)$, given any initial state ${x}\in{\mathcal{S}}^{*}$.